General Informations

Status-Key

Status/green.gif

Resolved

Status/orange.gif

Still working but with some errors

Status/red.gif

Pending


Webserver Outage

STATUS: Status/green.gif

2012-11-15: 14:30

Erroneous deletion of some apache configuration files led to outages of the webservers oenone and yosemite today between 13:30 and 14:15.

Email Service Outage

STATUS: Status/green.gif

2012-11-13: 07:30
VPP announces print service outage for about 15 min. . This should resolve the latest VPP service issues.


Email Service Outage

STATUS: Status/green.gif

2012-11-09: 10:05
Emails can not be sent/received. ISG.EE is working to resolve this problem as soon as possible.
2012-11-09: 12:20
We are currently still investigating what caused the outage of our mail server.
2012-11-09: 15:30

The amavis Daemon (a high-performance interface between mailer (MTA) and AVI content checker) stopped working because his temporary directories were removed. It is not clear what removed these directories. We are still investigating this but in the meantime the mailserver is up and running. NO EMAILS WERE LOST but the mails were sent/received with a 1-2 hours delay.


VPP Outage

STATUS: Status/green.gif

2012-11-02: 16:00
Jobs sent to VPP Printers don't get printed. We are investigating this problem together with VPP.
2012-11-05: 07:00
Services up and running again.


Complete Network Outage

STATUS: Status/green.gif

2012-10-25: 07:05 - 09:15
Complete network outage. Cause is still unclear but might be a side effect of the router hardware replacement due to a defect hardware announced yesterday.
2012-10-25: 09:15

The Informatik Dienste have posted a statement on their status page


Complete Network Outage

STATUS: Status/green.gif

2012-10-16: 20:20 - 20:40 / 22:10 - 23:20
Complete network outtage could be a side effect of a central router upgrade from informatik dienste. We are investigating it.
2012-10-17: 08:52

The informatik dienste have posted a statement on their status page


Outage Poster printer at ETZ J 66 / ETZSPEZ - 16. October 2012

STATUS: Status/green.gif

2012-10-16: 08:30 - 12:30
plotter HP6100 at ETZ J66 will be under maintenance and is not available during that time


Mailinglist Downtime (lists.ee.ethz.ch)

STATUS: Status/green.gif

2012-10-10: 09:00 - 17:00

Due to the migration of our mailing list software to sympa, we will take down the lists.ee.ethz.ch website. Mails sent to lists.ee.ethz.ch will put into a HOLD queue and delivered to the mailinglist after migration. So no mails get lost!

2012-10-10: 17:04
Mailinglists converted. Services up on running.


LDAP Migration

STATUS: Status/green.gif

2012-09-08: 10:00 - 17:00
All systems and services are not available during migration
2012-09-08: 17:30
All systems are back online


Network outage ETH

STATUS: Status/green.gif

2012-08-17: 16:30
Currently on ETH network there seem to be problems related on networking level. No information is available yet.
2012-08-17: 17:15
Systems are back online.


Power outage at ETH

STATUS: Status/green.gif

2012-06-28: 19:17
ETH had power outage affecting many services. D-ITET infrastructure was partially affected too.
2012-06-29: 08:00
We are currently working on resolving the outstanding issues and bringing back online services which are still down.
2012-06-29: 09:30

On ID Status website https://www1.ethz.ch/id/servicedesk/sysstat/index you can find now further information.

2012-06-29: 10:05
The cause was a transformator on fire in the main building, causing a power downtime in the computer centres from 19:30 to 22:00.


Server oenone crash

STATUS: Status/green.gif

2012-06-20: 00:00

oenone crashed. Affecting users of AUT, BIWI, COLLEGIUM, IKT and VAW.

2012-06-05: 00:30
oenone rebooted and is running again.

(!) We are still investigating what caused the crash and will report further information here.


oenone: hanging lockd affecting some User homes and webpages

STATUS: Status/green.gif

2012-06-05: 21:37

oenone is unresponsive. Affecting users of AUT, BIWI, COLLEGIUM, IKT and VAW.

2012-06-05: 00:00
oenone was rebooted and is running again.


Outage on 2/3 of alumni.ethz.ch mailservers

STATUS: Status/green.gif

2012-06-04: 16:00
Our mailserver can't deliver emails to alumni.ethz.ch addresses. Reason: The receiving Servers have temporary errors: "451 unable to verify user". Looks like something is misconfigured there.
2012-06-05: 10:00
Looks like all alumni servers hosted on tophost.ch have a problem. We have created a temporary transport map which delivers the mails to the genotec.ch (the only one which works) alumni mailserver. As long as you use our mailserver to send emails to alumni addresses they will be delivered immediately.


oenone: hanging lockd affecting some User homes and webpages

STATUS: Status/green.gif

2012-05-13: 22:50

oenone is unresponsive. Affecting users of AUT, BIWI, COLLEGIUM, IKT and VAW.

2012-05-14: 00:15
oenone was rebooted and is running again.


drwho: Main server outage affecting 64bit diskless Linux clients

STATUS: Status/green.gif

2012-05-08: 11:50
We currently experience some problems with one of our main server. All 64bit diskless clients are affected. We are working on the Problem. Furthermore some svn repositories might be affected.
2012-05-08: 13:15
The system is going back to normal but needs some time to fully recover.
2012-05-08: 13:45
The system should be back to normal. We still are working on some single hosts to recover them.


Maintenance work on D-ITET's central IT infrastructure

STATUS: Status/green.gif

2012-03-13: 21:30
All systems are back online.
2012-03-13: 19:00 - 22:00
To keep our systems up to date with the newest software and security releases, we need to update our servers on a regular basis. For this reason we are going to install latest Oracle patches on our main servers. The servers will be rebooted during this maintenance

To prevent data corruption/loss please do the following:

More details:


Maintenance downtime of Server behind ipp2vpp.ee.ethz.ch (printing/licenses/dhcp for selfmanaged hosts)

STATUS: Status/green.gif

On 2012-02-16 around 8:00AM we are scheduling a maintenance downtime of zaan (serving printing, license Server and DHCP Server for D-ITET). The downtime will be as short as possible. We plan to have it down for around 45 minutes at most.


oenone home server crash

STATUS: Status/green.gif

2012-02-13

During this night at around 22:40 oenone one of our home-servers crashed. Users with homes on oenone where affected, these are BIWI, VAW, Collegium Helveticum, Control, ISI, IKT. The system was up again at 00:40.

We are sorry for the caused inconvenience and we are investigating the problem.


Emergency reboot of ITET's server OENONE

STATUS: Status/green.gif

2012-01-31: 21:00

Emergency reboot of server OENONE due to a not responding storage area. User concerned: BIWI, VAW, Collegium Helveticum, Control, ISI, IKT, EEH

2012-01-31: 21:45
Server oenone is up again and all services are running


Emergency reboots of all Linux Clients and Servers

STATUS: Status/green.gif

2012-01-23: 15:00 PM
Due to a critical issue we were forced to reboot all affecting hosts We are sorry for the short notice and the inconvenience caused to you.
2012-01-23: 20:45 PM
All Clients rebooted.


CategoryEDUC

Status/Archive/2012 (last edited 2016-01-06 12:49:57 by alders)